OpenStudy (anonymous):

If f(x)=5^x find f(x+h)-f(x)/h -> f(x+h)-f(x) is all over h)

OpenStudy (anonymous):

\[\frac{ 5^{x+h}-5^{x} }{ h }\]\[\frac{ (x+h)\ln5-x \ln5 }{ \ln h }\rightarrow \frac{ h \ln 5 }{ \ln h }\]I think I'm wrong but maybe seeing this will give you an idea. Don't take this as fact.